Integrate cosecx.sec3x
 

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\\int \csc \left(x\right)\sec ^3\left(x\right)dx\:=\:\frac{\sec ^2\left(x\right)}{2}+\ln \left(\sin \left(x\right)\right)-\ln \left(\cos \left(x\right)\right)+C
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty